### v1

# GeoBenchX: Benchmarking LLMs for Multistep Geospatial Tasks

Abstract
In this paper, we establish a benchmark for evaluating large language models (LLMs) on multi-step geospatial tasks relevant to commercial GIS practitioners. We assess seven leading commercial LLMs (Sonnet 3.5 and 3.7, Haiku 3.5, Gemini 2.0, GPT-4o, GPT-4o mini, and o3-mini) using a simple tool-calling agent equipped with 23 geospatial functions. Our benchmark comprises tasks across four categories of increasing complexity, with both solvable and intentionally unsolvable tasks to test hallucination rejection. We develop an LLM-as-Judge evaluation framework to compare agent solutions against reference implementations. Results show Sonnet 3.5 and GPT-4o achieve the best overall performance, with Claude models excelling on solvable tasks while OpenAI models better identify unsolvable scenarios. We observe significant differences in token usage, with Anthropic models consuming substantially more tokens than competitors. Common errors include misunderstanding geometrical relationships, relying on outdated knowledge, and inefficient data manipulation. The resulting benchmark set, evaluation framework, and data generation pipeline are released as open-source resources, providing one more standardized method for ongoing evaluation of LLMs for GeoAI.
GeoBenchX: Benchmarking LLMs for Multistep Geospatial Tasks
Varvara Krechetova World Bank New York, USA                        Denis Kochedykov J.P.Morgan Chase New York, USA
1 Introduction
In this paper, we create a benchmark set and describe a simple LLM-as-Judge based methodology to evaluate the performance of several leading commercial large language models (LLMs) with function-calling abilities for solving a range of multi-step geospatial tasks. Geospatial tasks are distinct in that they require LLMs to have good factual geographical knowledge, understanding how to work with spatially defined data, georeferenced data, what a coordinate reference system is, understanding what are various types of spatial or geodata. We used a simple tool-calling agent to specifically assess the LLMs capabilities and have tested several models: Sonnet 3.5 and 3.7, Haiku 3.5, Gemini 2.0, GPT-4o, GPT-4o mini, o3-mini.
Our focus is on the commercial GIS practitioner requirements, which influence the choice of tasks, models and evaluation framework. Unlike other adjacent benchmarks, we focus on tasks that are closer to what GIS practitioners usually need for socio-economic analysis and mapping for reports, news articles, memos, and the like. Practitioners would use LLM equipped with GIS tools in a chat environment to get series of maps, numbers, and lists of locations. The tasks in the benchmark set reflect this – asking to make various types of map, to find geographical objects by some conditions, count objects within certain geographic boundaries, etc. Also, importantly for practitioners, we test the ability of LLMs to correctly recognize when the task cannot be solved with given tools and inform the user. This reduces the risk of hallucinated outputs. Next, we benchmark most current advanced commercial LLMs used in industry setting. One requirement of such industry usage is that the variability of outputs is not desired, so the models were tested at temperature 0. Lastly, the evaluation method we propose allows one to automate the ongoing evaluation which is increasingly important in practice with frequent models updates in commercial settings.
The benchmark set and code are publicly released at https://github.com/Solirinai/GeoBenchX.
The paper is structured as follows. First, we describe the task-solving agent and the set of tools/datasets. Second, we describe how the benchmark set of tasks was created and its composition. Third, we detail the approach to evaluation of the results of the task-solver agent using an LLM-as-Judge evaluator agent and testing the evaluator performance. Lastly, we compare the performance metrics of the solver agent based on different LLMs and the cost efficiency in terms of token usage.
2 Related works
Research in the field of GeoAI, a part of spatial data science at the intersection with Artificial Intelligence (AI) Janowicz et al. (2020), got a boost with recent advances in generative AI and the deployment of accessible LLMs. GeoAI researchers actively explore the capabilities of the available models in relation to geospatial data science, design their applications, and methods to improve their performance. The avenues explored include analysis of the ability of LLM (GPT-4) for spatial reasoning in travel networks, distance and elevation estimations, route and supply chain analysis Roberts et al. (2023); creating autonomous LLM-based GIS applications (LLM-Geo) Li and and (2023), GIS Copilot for GIS software Akinboyewa et al. (2024), ability LLMs to correctly understand and apply GIS toolsZhang et al. (2024), integrating LLM and cartography tools into MapGPT He et al. (2024), combining several LLMs to increase the efficiency of spatial data processing in the multi-agent framework for automated processing of vector data in shapefile format Lin et al. (2024), benefits of generative AI agents in smart mobilityXu et al. (2024), fine-tuning LLM to process natural language queries into executable codeMansourian and Oucheikh (2024). There is a benchmarking of LLM performance in spatial tasks in progress, with analysis of LLMs’ geospatial knowledgeBhandari et al. (2023); Renshaw et al. (2025), of GIS coding abilitiesGramacki et al. (2024), the recent study testing models on several groups of tasks that require geographic, GIS, and spatial knowledge, including changes in their performance in response to changes in prompts Liuchang Xu et al. (2025).
3 Agent, tools, and datasets
The experiment setup is based on the pre-built Langgraph ReAct agent LangChain AI (2024) equipped to call one of the selected LLMs with a list of attached tools. We intentionally choose a simple agent architecture to measure the impact of the LLMs on the final performance, not compensated and leveled by complex agent architectures.
The agent processed a given user task without Human-In-The-Loop involvement. Each task normally requires series of tool calls to solve and the agent does not always find the shortest solutions, so we allowed up to 25 iterations “LLM suggests tool call(s) – LLM receives the tool(s) response”. Less iterations were not enough for some complex tasks, while allowing more steps did not increase success rates. Experiments indicated that an increase in temperature leads to the agent being stuck looping over the same steps and to lower task completion rates, so, all LLMs with enabled temperature parameter were tested at temp. 0.
The agent system prompt sets the persona, briefly describes the setting, and the rules for solving the tasks (complete prompt in Appendix). The agent was instructed not to ask the user for clarifications and not to provide code to solve the task, but to respond with the resulting map, the names of the data sets used and explanations. The set of rules instructs when to reject the task and when to leave the solution empty (meaning that no tools are needed to answer user’s question), when to consider the data available (if the available data do not fully cover the geography asked in the task), and how to select columns for mapping data if the user did not specify (implied conditions frequent in production user questions).
The text of the task is supplied as the next message for the LLM Chat API as the user prompt.
Agent is provided with 23 tools (specifications are in the Appendix) that can read tabular data, geospatial data in vector or raster formats, perform operations on dataframes and geodataframes (merge, filter by categorical or numerical values), their columns, perform spatial operations on geodataframes and rasters (spatial joins, filtering by locations, creating buffers, generate contour lines from raster), visualize geodataframes, generate choropleth maps and heatmaps. The tools leverage GIS-specific libraries like GeoPandas den Bossche et al. (2024), Rasterio Gillies et al. (2013–), GDAL Rouault et al. (2025), Shapely Gillies et al. (2025), and common data analysis libraries like Matplotlib, Plotly, Pandas, and Numpy.
To distinguish the situations in which the agent stops solving the task because it decides that the task could not be solved with the provided data sets / tools, the agent was equipped with an additional tool rejecttask that simply returned the message that the task is unsolvable.
When a tool result included a dataframe or geodataframe, information about its length, columns, data types, and non-empty cells was included in the tool response message. Where dataframes or geodataframes needed to be passed between tools, they were stored in the data-store dictionary accessible to all tools.
For clarity, hereafter we refer to the set of tasks used to evaluate the agent as “benchmark set” and the statistical and geo data used by the tools to solve the tasks as “data sets”.
To make the experimentation consistent, we opted to store the snapshots of datasets in the local experimentation environment rather than fetching data from online APIs on every run. This allowed us to avoid variability of benchmarking due to possible data API connection issues, data API changes, and data changes behind the scenes.
Snapshots of data sets were organized as dictionaries with short descriptive titles as keys and filenames as values, one dictionary per type of data. When the tool designed to load the respective data was called, LLM had to choose the tool argument, the data set to load, from the list of the keys of the appropriate dictionary.
The data included: Statistical data: panel data in csv format (18 datasets): World Bank data on economic and population indicators, tuberculosis cases at subnational levels. Geometry data: vector data in shapefile format (21 datasets): polygon geometries (national or subnational units, lakes), points geometries (wildfire locations, urban places, Amtrak stations, earthquake occupancies, seaports, mineral extraction facilities, power stations), line geometries (rivers, railways). Raster data: (11 rasters): accumulated snow cover, flood extents, population data.
4 Benchmark set
To represent typical use by GIS professionals of a GIS assistant, tasks were single-sentence questions in natural language that might be asked by someone who needs a map, a number, a visualization for drafting an analytical research note, creating an action plan in international development, cross-checking information, etc. Some tasks were naturally ambiguous or missing information.
We created a set of 200+ tasks, grouped by type and complexity. Based on Wald’s method, this sample size corresponds approximately to the 5% margin of error for estimating a binary match rate with the 90% confidence. Each group included tasks that could be solved with the tools/datasets provided (solvable tasks), tasks that looked similar to solvable but could not be solved with the tools or datasets provided (unsolvable tasks). The groups of tasks are as follows:
Merge-visualize
(36 tasks): require to join tabular numeric data (panel data for one or multiple indicators, e.g. countries-by-years) with a geographic geometry data and make a choropleth or bivariate map. Example: Map the relationship between GDP per capita and electric power consumption per capita globally”. This is the most straightforward group, the tasks mostly require selection of correct datasets, and two (three in case of bivariate map) tool calls. Only 6 of these tasks were not solvable with given tools/data. Due to simplicity of the tasks, their content and solutions have limited variation; therefore, a limited number of them needed to have enough coverage.
Process-merge-visualize
(56 tasks): require some processing of either tabular or geometry data (filtering or operations on columns) and then creating a choropleth or bivariate map. Example: Compare rural population percentages in Western Asian countries”. The tasks in this group are more complicated, require the use of knowledge of the world geography and the use of a wider variety of tools. Half of the tasks, 28, are not solvable with given tools/data.
Spatial operations
(53 tasks): require spatial joins, creating buffers around selected geometries, checking overlaps between rasters, performing calculations on rasters, finding values from raster using geometry data or filtering geometries based on raster values, calculating distances. They might require a numerical answer, and the creation of actual maps might be optional. This group is more challenging than the previous two, as it requires the agent to work with different types of data set in one task, navigate the full set of available tools, and maintain the overall logic of the solution through all its steps. Example: How many people live within 1 km from a railway in Bangladesh?”. In this group, 23 of the tasks were not solvable.
Heatmaps, contour lines
(54 tasks): required performing spatial operations (selection, overlaps, etc.), understanding the contents of rasters, and making a heatmap or generating contour lines as one of the steps. Example: Create a heatmap of earthquake occurrences in regions with high population growth”. This was the most challenging set, as the agent had to figure out if the heatmaps and contour lines tools were actually able to solve the task and to make correct decisions on a complex set of arguments. In this group, 41 tasks ended up being unsolvable with available data/tools, which was a test for LLMs ability to recognize an unsolvable task and generate a rejection rather than a hallucinated answer.
4.1 Tasks creation
For each group, seed tasks were sourced from industry professionals and publications. Then more tasks were generated by using the Sonnet 3.5 model based on the seed tasks, the description of the tools, and the lists of available datasets. Then tasks were generated that should look similar to solvable tasks but could not be solved using the provided data and tools.
All the generated tasks were then manually reviewed to keep only high-quality tasks in the set.
4.2 Tasks annotation
Each of the 200+ tasks was annotated with one or multiple ground-truth reference solutions. To create the annotation, we have started by generating the draft reference solutions by the agent using GPT-4o or Sonnet 3.5 models. We then manually reviewed and rewrote every solution.
One reference solution could be (1) Sequence of tool calls with arguments; similar to a Python code snippet (2) Empty sequence [] of tool calls meaning that the task can be solved from the LLM’s world knowledge (3) A single special tool call reject_task() meaning that the task should be rejected as it cannot be solved with given tools/data.
A nonempty reference solution looked as follows for the evaluator:
#overall reference solution comment
#toolcall1 comment
toolcall1(arg11=..., arg12=..., ...)
#toolcall2 comment
toolcall2(arg21=..., arg12=..., ...)
Approximately 75% of the solvable tasks have reference solutions consisting of 4-6 steps (tool-calls) and  25% consist of more steps (up to 14 steps).
We used comments in annotations as a way to record information important for the evaluator agent to make a judgment about whether the candidate solution and the reference solution are semantically equivalent. For example, telling which variations in the solution or step are irrelevant or important.
One task sometimes allowed multiple solutions because tasks in practice are often ambiguous and multiple analytical approaches are possible for them. For example, task Compare the freshwater withdrawal between African countries with and without significant railway networks” can be solved: (1) by rejecting because it is not specific enough (user need to clarify what’s “signigicant” railway networks; (2) by making separate maps of railway networks and water withdrawals for African countries; (3) by combining the data and mapping bivariate map to show the relation; (4)by implying some threshold and computing statistical metrics for 2 groups of countries.
5 Evaluation approach
5.1 Matching-based evaluator
The 2 widely used evaluation approaches are execution-based and matching-based. However, comparing directly the outputs of the agent on geospatial analytical tasks would be quite challenging because the outputs contain multiple images/maps, tables, numbers, a narrative and sometime all of it as in the section 4.2 example. Annotation of such ground truth is not practical in this case, so we opted for matching-based approach. We used LLM-as-Judge with multiple LLM judges to compare the solutions generated by the agent (“candidate solutions”) and reference solutions for the same task. To leverage the LLMs ability to analyze code, we represented both reference and candidate solutions as snippets of Python code as in 4.2. We asked LLM judges to compare semantic equivalence of the solution; this took care of situations where reference and candidate solutions had different order of steps, different arguments for tool-calls or even took a different path to the same analytical outcome.
We also compared the lengths of the reference and candidate solutions, since the agent often takes some extra steps to figure out a path to answering the user question, and such extra steps are often costly.
For evaluation of the solutions, we used a panel of judges: Sonnet 3.5, GPT-4o, and Gemini 2.0. The evaluator prompt included: persona setting, description of all tools, scoring taxonomy, instructions on how evaluate solutions, examples of evaluated solutions, corresponding scores, and reasoning behind each score. The evaluator was instructed to analyze the candidate solution for the task and produce (1) reasoning for matching any of the reference solutions for the task and the candidate solution and (2) the matching score (‘2’: match, ‘0’: no match, ‘1’: partial match)
5.2 Evaluating the evaluator
We evaluated and tuned the evaluator itself on a set of 44 tasks randomly sampled with stratification of all tasks, respecting task groupings. We manually scored candidate solutions for these tasks, providing a matching score and reasoning for it. We then ran the evaluator on the same tasks candidate solutions and compared human annotator match scores, and the evaluator produced match scores.
Among the three models used as judges, the Sonnet 3.5 model matching scores were closest to human annotator scores. scoring (0.91 of the scores corresponded to manual scores), followed by GPT-4o (0.86) and Gemini 2.0 (0.82).
In the below results section, we use all the 3 above LLMs as judges and average their match frequency estimates.
6 LLMs evaluation results
The following models were tested in the solver agent: gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18, o3-mini-2025-01-31, gemini-2.0-flas%h-001, claude-3-5-sonnet-20241022, claude-3-5-haiku-20241022, claude-3-7-sonne%t-20250219. Each LLM was tested on the same set of tasks, and all models were tested at temperature 0 (except o3-mini, which does not have a temperature parameter).
The frequency of each of the match scores was calculated (1) separately for the solvable tasks (104 tasks) and the unsolvable tasks (98 tasks) (2) for each group of tasks according to 4.
To mark unsolvable tasks, the agent was provided with a rejecttask tool that it could call either upfront or after some steps if it reached the conclusion that the task cannot be solved with available datasets and tools.
To benchmark how efficient an LLM is at solving the task compared to the reference solution, we calculated the mean difference between the length (number of steps) of the reference solution (or average length of reference solutions if multiple) and the length of candidate solution. We express this difference relative to the length of the reference solution (so 0.5 means 50% more steps in the candidate solution than the reference).
The first observation is that models that are more successful in identifying unsolvable tasks tend to be less successful in dealing with solvable tasks and vice versa (Table 1, Table 2). Anthropic models perform better on solvable tasks, but prioritize giving any answer at all over rejecting tasks, Sonnet 3.5 is the most balanced from the three in this aspect. Gemini 2.0 performs similarly on both kinds of tasks. Although the three OpenAI models are better at rejecting tasks than solving them, among them, the GPT-4o is well balanced. Overall, on solvable tasks, Sonnet 3.5 turned out to be the best model, followed by Sonnet 3.7, GPT-4o, Gemini 2.0 and Haiku 3.5 (Table 2). We can also see that for unsolvable tasks, the candidate solutions are much longer, indicating LLMs trying to find a solution and eventually giving up. For solvable tasks, candidate solutions are about the same length as reference ones, with notable exceptions of Sonnet 3.7 that tends to produce longer solutions than needed and o3-mini that tends to reject tasks more frequently than other models. The detailed statistics for each group of tasks are in tables 3, 4, 5, 6.
Model No Partial Match $\Delta$ in
match match solution
length
Sonnet 3.7 0.84 0.02 0.14 6.47
Sonnet 3.5 0.61 0.02 0.37 2.51
Haiku 3.5 0.79 0.01 0.20 4.39
GPT-4o 0.38 0.03 0.59 2.45
GPT-4o mini 0.37 0.02 0.62 2.71
o3-mini 0.23 0.02 0.76 0.20
Gemini 2.0 0.54 - 0.46 1.83
Table 1: Success rate in identifying the unsolvable tasks (98 tasks from total of 202)
Model No Partial Match $\Delta$ in
match match solution
length
Sonnet 3.7 0.27 0.23 0.50 0.39
Sonnet 3.5 0.28 0.19 0.53 -0.05
Haiku 3.5 0.34 0.23 0.43 0.03
GPT-4o 0.41 0.15 0.44 -0.01
GPT-4o mini 0.70 0.13 0.16 -0.03
o3-mini 0.69 0.10 0.21 -0.59
Gemini 2.0 0.43 0.14 0.43 -0.03
Table 2: Success rate in solvable tasks (104 tasks from total of 202)
Model No Partial Match $\Delta$ in
match match solutions
length
Sonnet 3.7 0.23 0.11 0.66 0.91
Sonnet 3.5 0.17 0.15 0.72 0.29
Haiku 3.5 0.25 0.21 0.54 0.51
GPT-4o 0.24 0.13 0.63 0.31
GPT-4o mini 0.49 0.19 0.32 0.31
o3-mini 0.50 0.14 0.36 -0.39
Gemini 2.0 0.25 0.06 0.68 0.16
Table 3: Success rate in solving ‘Merge-visualize’ tasks (36 tasks from total of 202)
Model No Partial Match $\Delta$ in
match match solution
length
Sonnet 3.7 0.50 0.14 0.36 2.37
Sonnet 3.5 0.30 0.10 0.60 0.40
Haiku 3.5 0.59 0.09 0.33 2.11
GPT-4o 0.26 0.09 0.65 0.29
GPT-4o mini 0.54 0.07 0.39 0.60
o3-mini 0.48 0.05 0.46 -0.38
Gemini 2.0 0.35 0.12 0.53 0.17
Table 4: Success rate in solving ‘Process-merge-visualize’ tasks (56 tasks from total of 202)
Model No Partial Match $\Delta$ in
match match solution
length
Sonnet 3.7 0.55 0.14 0.31 3.42
Sonnet 3.5 0.48 0.12 0.40 1.73
Haiku 3.5 0.50 0.17 0.33 2.49
GPT-4o 0.50 0.08 0.42 1.95
GPT-4o mini 0.69 0.05 0.26 2.44
o3-mini 0.55 0.07 0.39 -0.3
Gemini 2.0 0.63 0.07 0.30 2.07
Table 5: Success rate in solving ‘Spatial operations’ tasks (53 tasks from total of 202)
Model No Partial Match $\Delta$ in
match match solution
length
Sonnet 3.7 0.81 0.13 0.06 6.05
Sonnet 3.5 0.76 0.10 0.13 2.15
Haiku 3.5 0.82 0.06 0.12 3.06
GPT-4o 0.56 0.08 0.36 2.00
GPT-4o mini 0.44 0.03 0.53 1.62
o3-mini 0.39 - 0.61 0.17
Gemini 2.0 0.64 0.04 0.32 0.91
Table 6: Success rate in solving ‘Heatmaps, contour lines’ tasks (54 tasks from total of 202)
Model Input Output
tokens (mln) tokens (k)
Sonnet 3.7 24 269
Sonnet 3.5 12 157
Haiku 3.5 15 203
GPT-4o 7.5 57
GPT-4o mini 8 55
o3-mini 3.5 561
Gemini 2.0 7 40
Table 7: Tokens used for generation of solutions for the whole set
6.1 Error Analysis
Looking at cases where candidate solutions did not match reference, the main causes/types of errors were:
(1) Poor understanding of geometrical relations: mistaking centroids for full polygons, attempting to get centroids of lines, and not understanding that centroids can be outside object boundaries, etc, etc. (2) Outdated world knowledge – good for established international groupings but fails with recent changes (e.g., BRICS expansion in 2024) or ambiguous groupings (Upper/Lower Nile countries). (3) Over-reliance on world knowledge - using internally known spelling of names rather than dataset values for filtering, despite having tools to check column values. (4) Inefficient data manipulation – merging without pre-filtering (e.g., merging with all USA counties instead of first filtering to specific state). (5) Skipping geographic filtering steps – keeping continent-level data when only country-level was needed. (6) Attempting to solve with partial data, trying to be helpful by using whatever geography was available instead of rejecting tasks with insufficient data coverage. (7) Modifying tasks to make them solvable rather than solving the original problem.
7 Conclusions
We have created a benchmark dataset for multistep geo-spatial task completion, introduced an evaluation framework, and evaluated the leading mainstream commercial LLMs on these tasks. The task profile and the evaluation approach reflect the practical requirements of GIS professionals and the use of LLM in a GIS-assistant mode.
Experiments indicate that commercial LLMs with function calling can successfully solve geospatial tasks even with casual, ambiguous instructions. Performance varies by task complexity, spatial thinking requirements, and model helpfulness bias.
Sonnet 3.5 and GPT-4o showed the best overall results and a balance between identifying unsolvable tasks and solving solvable ones. Anthropic models performed well on easier tasks, but their preference for providing solutions rather than rejecting tasks, high token usage, and price may disadvantage them for geospatial tasks. o3-mini and GPT-4o-mini overall had lower success rates, especially in easy tasks, though they excel at identifying unsolvable tasks and are affordable.
Limitations
We did not tune prompts individually for each LLM; we evaluate only solution algorithm not the final response (for rejects and final numbers); the benchmark set size is moderate; we didn’t evaluate the variance of the responses (only the binomial confidence bounds).
Appendix A Prompts for solver agent
The system prompt is in Figure 1, the rules prompt is in Figure 2.
System Prompt You are a geographer who answers the questions with maps whenever possible. You do have a list of python functions that help you to load statistical data, geospatial data, rasters, merge dataframes, filter dataframes, get uniques values from columns, to plot a map, make spatial operations on raster and vector data. You do not provide code to generate a map, you provide names of the datasets you used; the statistical dataset, the geospatial dataset and the resulting map with legend and needed explanations. Today is {current_date}.
Figure 1: System prompt for solver agent.
Rules Prompt While solving the tasks, please, follow the next rules: - If a task is not a geospatial task, if it does not require calling provided tools and use provided datasets to solve, proceed with reply, no explanations needed, no need to call reject_task’. - If the task is a geospatial task, however, this task can NOT be solved with either available datasets or tools, call the tool reject_task’. - If user did not specify the date of the data to be plotted on the map, you map the latest available period that has non empty data for more than 70% of the objects in the datasets (example: if a question is about countries of Africa, you map the latest year that has data for more than 70% of countries in Africa). - For bivariate maps select the data for the same year for both variables. For instance, if for variable 1 the latest data are for 2023 and for variable 2 the latest data are for 2014, select data for 2014 for both variable 1 and variable 2. If the common year for which both datasets have data is 2012, take the data for 2012. - If available dataset or datasets cover only part of the requested geography, consider the data unavailable. - Do not ask questions to the user, proceed with solving the task till a result is achieved. When the information is missing try to infer the reasonable values.
Figure 2: Rules prompt for solver agent.
Appendix B GeoSpatial Tools List
loaddata
Loads statistical data from the catalog into a Pandas DataFrame.
  • dataset: Name of the dataset to load
  • outputdataframename: Name for the new DataFrame
loadgeodata
Loads vector geospatial data from the catalog into a GeoPandas GeoDataFrame.
  • geodataset: Name of the geodataset to load
  • outputgeodataframename: Name for the new GeoDataFrame
getrasterpath
Constructs path to raster data from the catalog.
  • rasterdataset: Name of the raster dataset to access
getrasterdescription
Gets description of a raster dataset including metadata and statistics.
  • rasterpath: Path to the raster file
  • outputvariablename: Optional name for storing results
analyzerasteroverlap
Analyzes overlap between two rasters and calculates statistics.
  • raster1path: Path to first raster file
  • raster2path: Path to second raster file
  • outputvariablename: Name for storing results
  • resamplingmethod1/2: Resampling methods
  • plotresult: Whether to display visualization
getvaluesfromrasterwithgeometries
Masks a raster using vector geometries and calculates statistics.
  • rasterpath: Path to the raster file
  • geodataframename: Name of GeoDataFrame with geometries
  • outputvariablename: Name for storing results
  • plotresult: Whether to display visualization
mergedataframes
Merges statistical and geospatial dataframes using key columns.
  • dataframename: Name of DataFrame with statistical data
  • geodataframename: Name of GeoDataFrame with spatial data
  • statkey: Column name in statistical data for joining
  • geokey: Column name in spatial data for joining
  • outputdataframename: Name for the merged DataFrame
getuniquevalues
Gets unique values from a column in a DataFrame/GeoDataFrame.
  • dataframename: Name of DataFrame to analyze
  • column: Column name to get unique values from
filtercategorical
Filters DataFrame/GeoDataFrame by categorical values.
  • dataframename: Name of DataFrame to filter
  • filters: Dictionary of column names and values
  • outputdataframename: Name for filtered DataFrame
filternumerical
Filters DataFrame/GeoDataFrame using numerical conditions.
  • dataframename: Name of DataFrame to filter
  • conditions: Query string with numerical conditions
  • outputdataframename: Name for filtered DataFrame
calculatecolumnstatistics
Calculates summary statistics for a numerical column.
  • dataframename: Name of DataFrame to analyze
  • columnname: Column to analyze
  • outputvariablename: Name for storing results
  • includequantiles: Whether to include standard quantiles
  • additionalquantiles: Additional quantiles to calculate
createbuffer
Creates buffer zones around geometries in a GeoDataFrame.
  • geodataframename: Name of source GeoDataFrame
  • buffersize: Size of the buffer in meters
  • outputgeodataframename: Name for result GeoDataFrame
makechoroplethmap
Creates a choropleth map visualization from a GeoDataFrame.
  • dataframename: Name of GeoDataFrame with map data
  • mappingkey: Column to visualize
  • legendtext: Legend title
  • colormap: Color scheme for the map
filterpointsbyrastervalues
Samples raster values at point locations and filters by threshold.
  • rasterpath: Path to the raster file
  • pointsgeodataframename: Name of GeoDataFrame with points
  • valuecolumn: Column to store raster values
  • outputgeodataframename: Name for filtered GeoDataFrame
  • filtertype: Type of threshold filter
  • threshold1: First threshold value
  • threshold2: Optional second threshold value
  • plotresult: Whether to display visualization
selectfeaturesbyspatialrelationship
Selects features based on spatial relationships.
  • featuresgeodataframename: GeoDataFrame to select from
  • referencegeodataframename: GeoDataFrame to select by
  • spatialpredicates: List of spatial relationships
  • outputgeodataframename: Name for result GeoDataFrame
  • plotresult: Whether to display visualization
calculatelinelengths
Calculates lengths of line features in kilometers.
  • geodataframename: Name of GeoDataFrame with line features
  • outputvariablename: Name for storing results
calculatecolumns
Performs operations between columns of DataFrames.
  • dataframe1name: Name of first DataFrame
  • column1name: Column from first DataFrame
  • dataframe2name: Name of second DataFrame
  • column2name: Column from second DataFrame
  • operation: Mathematical operation to perform
  • outputcolumnname: Name for results column
  • outputdataframename: Name for result DataFrame
scalecolumnbyvalue
Performs operations between a column and a value.
  • dataframename: Name of DataFrame to modify
  • columnname: Column to scale
  • operation: Mathematical operation
  • value: Numerical value for scaling
  • outputcolumnname: Name for results column
  • outputdataframename: Name for result DataFrame
makeheatmap
Creates an interactive heatmap from point data.
  • geodataframename: Name of GeoDataFrame with points
  • valuecolumn: Column for intensity values
  • outputhtmlname: Optional name for HTML output
  • centerlat/lon: Map center coordinates
  • zoomlevel: Initial zoom level
  • radius: Radius of influence for each point
  • mapstyle: Base map style
  • width/height: Map dimensions
visualizegeographies
Displays multiple GeoDataFrames as map layers.
  • geodataframenames: List of GeoDataFrame names
  • layerstyles: Optional styling for each layer
  • basemapstyle: Style of the basemap
  • title: Map title
  • addlegend: Whether to add a legend
getcentroids
Calculates centroids of polygon features.
  • geodataframename: Name of GeoDataFrame with polygons
  • outputgeodataframename: Name for centroid GeoDataFrame
plotcontourlines
Creates contour lines from raster data.
  • rasterpath: Path to the raster file
  • outputgeodataframename: Name for contour GeoDataFrame
  • interval: Interval between contour lines
  • min/maxvalue: Range for contours
  • plotresult: Whether to display plot
  • title: Plot title
generatecontoursdisplay
Creates contour lines using GDAL algorithm.
  • rasterpath: Path to the raster file
  • outputfilename: Name for output shapefile
  • contourinterval: Interval between contour lines
  • columntitle: Name for values column
  • nodataval: Value for no-data pixels
  • outputgeodataframename: Name for contour GeoDataFrame
  • min/maxvalue: Range for contours
  • plotresult: Whether to display visualization
makebivariatemap
Creates a bivariate choropleth map showing relationship between two variables.
  • dataframename: Name of GeoDataFrame with map data
  • var1/var2: Variable column names
  • var1name/var2name: Display names for variables
  • title: Map title
  • outputvariablename: Name for storing results
rejecttask
Returns a standardized message when task cannot be solved.

### v2

@@
-GeoBenchX: Benchmarking LLMs for Multistep Geospatial Tasks
-Abstract
-In this paper, we establish a benchmark for evaluating large language models (LLMs) on multi-step geospatial tasks relevant to commercial GIS practitioners.
-We assess seven leading commercial LLMs (Sonnet 3.5 and 3.7, Haiku 3.5, Gemini 2.0, GPT-4o, GPT-4o mini, and o3-mini) using a simple tool-calling agent equipped with 23 geospatial functions.
-Our benchmark comprises tasks across four categories of increasing complexity, with both solvable and intentionally unsolvable tasks to test hallucination rejection.
-We develop an LLM-as-Judge evaluation framework to compare agent solutions against reference implementations.
-Results show Sonnet 3.5 and GPT-4o achieve the best overall performance, with Claude models excelling on solvable tasks while OpenAI models better identify unsolvable scenarios.
-We observe significant differences in token usage, with Anthropic models consuming substantially more tokens than competitors.
+GeoBenchX: Benchmarking LLMs in Agent Solving Multistep Geospatial Tasks
+Abstract.
+This paper establishes a benchmark for evaluating tool-calling capabilities of large language models (LLMs) on multi-step geospatial tasks relevant to commercial GIS practitioners.
+We assess eight commercial LLMs (Claude Sonnet 3.5 and 4, Claude Haiku 3.5, Gemini 2.0 Flash, Gemini 2.5 Pro Preview, GPT-4o, GPT-4.1 and o4-mini) using a simple tool-calling agent equipped with 23 geospatial functions.
+Our benchmark comprises tasks in four categories of increasing complexity, with both solvable and intentionally unsolvable tasks to test rejection accuracy.
+We develop a LLM-as-Judge evaluation framework to compare agent solutions against reference solutions.
+Results show o4-mini and Claude 3.5 Sonnet achieve the best overall performance, OpenAI’s GPT-4.1, GPT-4o and Google’s Gemini 2.5 Pro Preview do not fall far behind, but the last two are more efficient in identifying unsolvable tasks.
+Claude Sonnet 4, due its preference to provide any solution rather than reject a task, proved to be less accurate.
+We observe significant differences in token usage, with Anthropic models consuming more tokens than competitors.
 Common errors include misunderstanding geometrical relationships, relying on outdated knowledge, and inefficient data manipulation.
-The resulting benchmark set, evaluation framework, and data generation pipeline are released as open-source resources, providing one more standardized method for ongoing evaluation of LLMs for GeoAI.
-GeoBenchX: Benchmarking LLMs for Multistep Geospatial Tasks
-Varvara Krechetova World Bank New York, USA                        Denis Kochedykov J.P.Morgan Chase New York, USA
-1 Introduction
-In this paper, we create a benchmark set and describe a simple LLM-as-Judge based methodology to evaluate the performance of several leading commercial large language models (LLMs) with function-calling abilities for solving a range of multi-step geospatial tasks.
-Geospatial tasks are distinct in that they require LLMs to have good factual geographical knowledge, understanding how to work with spatially defined data, georeferenced data, what a coordinate reference system is, understanding what are various types of spatial or geodata.
-We used a simple tool-calling agent to specifically assess the LLMs capabilities and have tested several models: Sonnet 3.5 and 3.7, Haiku 3.5, Gemini 2.0, GPT-4o, GPT-4o mini, o3-mini.
-Our focus is on the commercial GIS practitioner requirements, which influence the choice of tasks, models and evaluation framework.
-Unlike other adjacent benchmarks, we focus on tasks that are closer to what GIS practitioners usually need for socio-economic analysis and mapping for reports, news articles, memos, and the like.
-Practitioners would use LLM equipped with GIS tools in a chat environment to get series of maps, numbers, and lists of locations.
-The tasks in the benchmark set reflect this – asking to make various types of map, to find geographical objects by some conditions, count objects within certain geographic boundaries, etc.
-Also, importantly for practitioners, we test the ability of LLMs to correctly recognize when the task cannot be solved with given tools and inform the user.
-This reduces the risk of hallucinated outputs.
-Next, we benchmark most current advanced commercial LLMs used in industry setting.
-One requirement of such industry usage is that the variability of outputs is not desired, so the models were tested at temperature 0.
-Lastly, the evaluation method we propose allows one to automate the ongoing evaluation which is increasingly important in practice with frequent models updates in commercial settings.
-The benchmark set and code are publicly released at https://github.com/Solirinai/GeoBenchX.
+The resulting benchmark set, evaluation framework, and data generation pipeline are released as open-source resources  [^1] Available at https://github.com/Solirinai/GeoBenchX, providing one more standardized method for the ongoing evaluation of LLMs for GeoAI.
+Large Language Models (LLMs), Geospatial Analysis, Function Calling, LLM-as-Judge Evaluation, GIS Tools, Benchmarks, Spatial Reasoning, Tool-based Agents
+journalyear: 2025copyright: acmlicensedconference: The 1st ACM SIGSPATIAL International Workshop on Generative and Agentic AI for Multi-Modality Space-Time Intelligence; November 3–6, 2025; Minneapolis, MN, USAbooktitle: The 1st ACM SIGSPATIAL International Workshop on Generative and Agentic AI for Multi-Modality Space-Time Intelligence (GeoGenAgent ’25), November 3–6, 2025, Minneapolis, MN, USAdoi: 10.1145/3764915.3770721isbn: /2025/11ccs: Computing methodologies Artificial intelligenceccs: General and reference General conference proceedingsccs: General and reference Evaluationccs: Computing methodologies Intelligent agentsccs: Computing methodologies Multi-agent systemsccs: Computing methodologies Information extraction
+1. Introduction
+In this paper, we create a benchmark set and describe a simple LLM-as-Judge based methodology to compare the tool-calling performance of several leading commercial large language models (LLMs) with function-calling abilities for solving a range of multi-step geospatial tasks.
+Geospatial tasks are distinct in that they require LLMs to have good factual geographical knowledge, understanding how to work with spatially defined data, what are coordinate reference systems, and various types of spatial data.
+We used a simple tool-calling agent to specifically assess the LLMs capabilities and have tested several models: Anthropic’s Claude Sonnet 3.5, 4, Claude Haiku 3.5; Google’s Gemini 2.0 Flash, Gemini Pro 2.5 Preview; OpenAI’s GPT-4o, GPT-4.1, o3-mini and o4-mini.
+Our focus is on the commercial GIS practitioner requirements, which influences the choice of tasks, models and evaluation framework.
+Unlike other adjacent benchmarks, we focus on tasks needed for socio-economic analysis and mapping for reports, policy papers, news articles, and the like: tasking agent to make various types of maps, to find geographical objects by some conditions, to count objects within certain geographic boundaries.
+For these tasks, practitioners would use LLM equipped with GIS tools in a chat environment.
+Also, importantly for practitioners, we test the ability of LLMs to correctly recognize when the task cannot be solved with given tools and inform the user to reduce the risk of hallucinated outputs.
+Next, we selected the most current advanced commercial LLMs used in the industry setting for benchmarking.
+Lastly, the evaluation method we use allows for automation of the ongoing evaluation which is increasingly important with frequent models updates in commercial settings.
 The paper is structured as follows.
-First, we describe the task-solving agent and the set of tools/datasets.
+First, we describe the task-solving agent and the set of tools.
 Second, we describe how the benchmark set of tasks was created and its composition.
-Third, we detail the approach to evaluation of the results of the task-solver agent using an LLM-as-Judge evaluator agent and testing the evaluator performance.
-Lastly, we compare the performance metrics of the solver agent based on different LLMs and the cost efficiency in terms of token usage.
-2 Related works
-Research in the field of GeoAI, a part of spatial data science at the intersection with Artificial Intelligence (AI) Janowicz et al.
-(2020), got a boost with recent advances in generative AI and the deployment of accessible LLMs.
+Third, we detail the approach to evaluation of the results of the task-solver agent using a panel of LLM-as-Judge evaluators and testing the evaluator performance itself.
+Finally, we compare the performance metrics of the solver agent based on different LLMs and the cost efficiency in terms of token usage.
+2. Related works
+Research in the field of GeoAI, a part of spatial data science at the intersection with Artificial Intelligence (AI) (Janowicz et al., 2020), got a boost with recent advances in generative AI and the deployment of accessible LLMs.
 GeoAI researchers actively explore the capabilities of the available models in relation to geospatial data science, design their applications, and methods to improve their performance.
-The avenues explored include analysis of the ability of LLM (GPT-4) for spatial reasoning in travel networks, distance and elevation estimations, route and supply chain analysis Roberts et al.
-(2023); creating autonomous LLM-based GIS applications (LLM-Geo) Li and and (2023), GIS Copilot for GIS software Akinboyewa et al.
-(2024), ability LLMs to correctly understand and apply GIS toolsZhang et al.
-(2024), integrating LLM and cartography tools into MapGPT He et al.
-(2024), combining several LLMs to increase the efficiency of spatial data processing in the multi-agent framework for automated processing of vector data in shapefile format Lin et al.
-(2024), benefits of generative AI agents in smart mobilityXu et al.
-(2024), fine-tuning LLM to process natural language queries into executable codeMansourian and Oucheikh (2024).
-There is a benchmarking of LLM performance in spatial tasks in progress, with analysis of LLMs’ geospatial knowledgeBhandari et al.
-(2023); Renshaw et al.
-(2025), of GIS coding abilitiesGramacki et al.
-(2024), the recent study testing models on several groups of tasks that require geographic, GIS, and spatial knowledge, including changes in their performance in response to changes in prompts Liuchang Xu et al.
-(2025).
-3 Agent, tools, and datasets
-The experiment setup is based on the pre-built Langgraph ReAct agent LangChain AI (2024) equipped to call one of the selected LLMs with a list of attached tools.
-We intentionally choose a simple agent architecture to measure the impact of the LLMs on the final performance, not compensated and leveled by complex agent architectures.
+The avenues explored include analysis of the ability of LLM (GPT-4) for spatial reasoning in travel networks, distance and elevation estimations, route and supply chain analysis (Roberts et al., 2023); integration of geography knowledge in pre-trained language models (Ramrakhiyani et al., 2025), creating autonomous LLM-based GIS applications (LLM-Geo) (Li and and, 2023), GIS Copilot for GIS software (Akinboyewa et al., 2024), integrating LLM and cartography tools into MapGPT (He et al., 2024), testing LLMs in styling maps using reference visuals (Wang et al., 2025), combining several LLMs to increase the efficiency of spatial data processing in the multi-agent framework for automated processing of vector data in shapefile format (Lin et al., 2024), benefits of generative AI agents in smart mobility(Xu et al., 2024), fine-tuning LLM to process natural language queries into executable code(Mansourian and Oucheikh, 2024).
+There is a benchmarking of LLM performance in spatial tasks in progress, with analysis of LLMs’ geospatial knowledge(Bhandari et al., 2023; Renshaw et al., 2025), of GIS coding abilities(Gramacki et al., 2024), the recent study testing models on several groups of tasks that require geographic, GIS, and spatial knowledge, including changes in their performance in response to changes in prompts (Xu et al., 2025).
+3. Experiment setup
+3.1. Agent
+The experiment setup is based on the pre-built Langgraph ReAct agent (LangChain AI, 2024) equipped to call one of the selected LLMs with a list of tools (functions) to solve tasks.
+We intentionally choose a simple agent architecture to measure the contribution of the LLMs themselves into the final performance, not compensated and leveled by complex agent architectures.
+Typical practical requirements are that the variability in outputs is undesirable, but consistency in responses is critical, so the models were tested at temperature 0, which reduces variability and increases consistency of outputs (OpenAI, 2025; Google Cloud, 2025).
 The agent processed a given user task without Human-In-The-Loop involvement.
-Each task normally requires series of tool calls to solve and the agent does not always find the shortest solutions, so we allowed up to 25 iterations “LLM suggests tool call(s) – LLM receives the tool(s) response”.
-Less iterations were not enough for some complex tasks, while allowing more steps did not increase success rates.
-Experiments indicated that an increase in temperature leads to the agent being stuck looping over the same steps and to lower task completion rates, so, all LLMs with enabled temperature parameter were tested at temp.
-0.
-The agent system prompt sets the persona, briefly describes the setting, and the rules for solving the tasks (complete prompt in Appendix).
-The agent was instructed not to ask the user for clarifications and not to provide code to solve the task, but to respond with the resulting map, the names of the data sets used and explanations.
-The set of rules instructs when to reject the task and when to leave the solution empty (meaning that no tools are needed to answer user’s question), when to consider the data available (if the available data do not fully cover the geography asked in the task), and how to select columns for mapping data if the user did not specify (implied conditions frequent in production user questions).
-The text of the task is supplied as the next message for the LLM Chat API as the user prompt.
-Agent is provided with 23 tools (specifications are in the Appendix) that can read tabular data, geospatial data in vector or raster formats, perform operations on dataframes and geodataframes (merge, filter by categorical or numerical values), their columns, perform spatial operations on geodataframes and rasters (spatial joins, filtering by locations, creating buffers, generate contour lines from raster), visualize geodataframes, generate choropleth maps and heatmaps.
-The tools leverage GIS-specific libraries like GeoPandas den Bossche et al.
-(2024), Rasterio Gillies et al.
-(2013–), GDAL Rouault et al.
-(2025), Shapely Gillies et al.
-(2025), and common data analysis libraries like Matplotlib, Plotly, Pandas, and Numpy.
-To distinguish the situations in which the agent stops solving the task because it decides that the task could not be solved with the provided data sets / tools, the agent was equipped with an additional tool rejecttask that simply returned the message that the task is unsolvable.
+Each task normally requires series of tool calls to solve and the agent does not always find the shortest solutions, so we allowed up to 25 iterations of “LLM suggests a tool call –¿ LLM receives the tool response”.
+This number was selected after initial trials of the agent that showed that fewer iterations were not enough for some complex tasks, while allowing more steps did not increase the success rates (accuracy).
+The agent is supplied with a simple system prompt that sets the persona, briefly describes the setting, and the rules for solving the tasks.
+The agent is explicitly instructed not to ask the user for clarifications and not to provide code to solve the task, but to reply with the resulting map, names of the datasets used, and explanations.
+The set of rules contains instructions on when to reject the task and when to leave the solution empty (meaning that no tools are needed to answer user’s question), when to consider the data unavailable (if the available data do not fully cover the geography asked in the task), and how to select columns for mapping data if the user did not specify (implied conditions frequent in production user questions).
+The text of the task is supplied as the next message for the LLM’s Chat API as the user prompt.
+3.2. Agent tools
+To enable the agent to solve the tasks, it is provided with 23 tools to implement the necessary data manipulations, calculations, and visualizations or to report a task unsolvable.
+Tools can read tabular data, geospatial data in vector or raster formats, perform operations on dataframes and geodataframes (merge, filter by categorical or numerical values), their columns, perform spatial operations on geodataframes and rasters (spatial joins, filtering by locations, creating buffers, generate contour lines from raster), visualize geodataframes, generate choropleth maps and heatmaps.
+The tools leverage GIS-specific libraries like GeoPandas (den Bossche et al., 2024), Rasterio (Gillies et al., 13), GDAL (Rouault et al., 2025), Shapely (Gillies et al., 2025), and common data analysis libraries like Matplotlib, Plotly, Pandas, and Numpy.
+To distinguish the situations in which the agent stops solving the task because it thinks that the task could not be solved with the datasets / tools provided, the agent was equipped with an additional tool rejecttask that simply returned the message that the task is unsolvable.
+The outputs of the tool functions were formatted as natural part of the tool response message.
 When a tool result included a dataframe or geodataframe, information about its length, columns, data types, and non-empty cells was included in the tool response message.
@@
 For clarity, hereafter we refer to the set of tasks used to evaluate the agent as “benchmark set” and the statistical and geodata used by the tools to solve the tasks as “datasets”.
-To make the experimentation consistent, we opted to store the snapshots of datasets in the local experimentation environment rather than fetching data from online APIs on every run.
-This allowed us to avoid variability of benchmarking due to possible data API connection issues, data API changes, and data changes behind the scenes.
-Snapshots of data sets were organized as dictionaries with short descriptive titles as keys and filenames as values, one dictionary per type of data.
-When the tool designed to load the respective data was called, LLM had to choose the tool argument, the data set to load, from the list of the keys of the appropriate dictionary.
-The data included: Statistical data: panel data in csv format (18 datasets): World Bank data on economic and population indicators, tuberculosis cases at subnational levels.
-Geometry data: vector data in shapefile format (21 datasets): polygon geometries (national or subnational units, lakes), points geometries (wildfire locations, urban places, Amtrak stations, earthquake occupancies, seaports, mineral extraction facilities, power stations), line geometries (rivers, railways).
-Raster data: (11 rasters): accumulated snow cover, flood extents, population data.
-4 Benchmark set
-To represent typical use by GIS professionals of a GIS assistant, tasks were single-sentence questions in natural language that might be asked by someone who needs a map, a number, a visualization for drafting an analytical research note, creating an action plan in international development, cross-checking information, etc.
-Some tasks were naturally ambiguous or missing information.
+To avoid variability of benchmarking due to possible data API connection issues, data API changes, and data changes behind the scenes, we opt to store the snapshots of datasets in the local experimentation environment rather than fetching data from online APIs every time.
+For each task, the conversation history between LLM and tools along with any plotted outputs is saved as html file making the final LLM response, maps and intermediate outputs available for debugging.
+3.3. Datasets
+The datasets were used as-is after downloading from online APIs, except for two modifications made to ease development: country names in the vector geodataset were harmonized with statistical datasets for easier visual identification of filtering issues, and the NoData value in snow cover rasters was unified.
+These changes were not required for the benchmarking itself.
+The data included:
+Statistical data: panel data in csv format (18 datasets): data on economic and population indicators, data on CO and greenhouse gas emissions, tuberculosis cases at subnational levels.
+Geometry data: vector data in shapefile format (21 datasets): polygon geometries (national or subnational units, lakes), points geometries (wildfire locations, urban places, Amtrak stations, earthquake occurrences, seaports, mineral extraction facilities, power stations at global or national level), line geometries (rivers, railways continent or national extent).
+Raster data: (11 rasters): accumulated snow cover, flood extents, population data at national level.
+4. Benchmark set
 We created a set of 200+ tasks, grouped by type and complexity.
-Based on Wald’s method, this sample size corresponds approximately to the 5% margin of error for estimating a binary match rate with the 90% confidence.
-Each group included tasks that could be solved with the tools/datasets provided (solvable tasks), tasks that looked similar to solvable but could not be solved with the tools or datasets provided (unsolvable tasks).
-The groups of tasks are as follows:
-Merge-visualize
-(36 tasks): require to join tabular numeric data (panel data for one or multiple indicators, e.g.
+For estimating a binary match rate at the level of match we see in results section (6) – this sample size corresponds approximately to 5% margin of error with 90% confidence according to Wald method.
+Each group included tasks that could be solved with the tools and datasets provided (solvable tasks), tasks that looked similar but could not be solved with the tools or datasets provided (unsolvable tasks).
+For debugging purposes, we also added a few tasks that did not require use of the provided tools or datasets to solve.
+4.1. Groups
+Depending on the complexity of data processing or geoprocessing, the task groups are as follows:
+Merge - visualize (36 tasks): require to join tabular numeric data (panel data for one or multiple indicators, e.g.
 countries-by-years) with a geographic geometry data and make a choropleth or bivariate map.
 Example: Map the relationship between GDP per capita and electric power consumption per capita globally”.
-This is the most straightforward group, the tasks mostly require selection of correct datasets, and two (three in case of bivariate map) tool calls.
+This is the most straightforward group, the tasks mostly require two-three tool calls.
 Only 6 of these tasks were not solvable with given tools/data.
-Due to simplicity of the tasks, their content and solutions have limited variation; therefore, a limited number of them needed to have enough coverage.
-Process-merge-visualize
-(56 tasks): require some processing of either tabular or geometry data (filtering or operations on columns) and then creating a choropleth or bivariate map.
+Due to simplicity of the tasks, there is limited variation in their content and solutions and a smaller number provided enough coverage.
+Process - merge - visualize (56 tasks): require some processing of either tabular or geometry data (filtering or operations on columns) and then creating a choropleth or bivariate map.
 Example: Compare rural population percentages in Western Asian countries”.
 The tasks in this group are more complicated, require the use of knowledge of the world geography and the use of a wider variety of tools.
-Half of the tasks, 28, are not solvable with given tools/data.
-Spatial operations
-(53 tasks): require spatial joins, creating buffers around selected geometries, checking overlaps between rasters, performing calculations on rasters, finding values from raster using geometry data or filtering geometries based on raster values, calculating distances.
+About half of the tasks, 24, are not solvable with given tools/data.
+Spatial operations (53 tasks): require spatial joins, creating buffers around selected geometries, checking overlaps between rasters, performing calculations on rasters, finding values from raster using geometry data or filtering geometries based on raster values, calculating distances.
 They might require a numerical answer, and the creation of actual maps might be optional.
-This group is more challenging than the previous two, as it requires the agent to work with different types of data set in one task, navigate the full set of available tools, and maintain the overall logic of the solution through all its steps.
+This group is more challenging than the previous two, as it requires the agent to work with different types of datasets in one task, navigate the full set of available tools, and maintain the overall logic of the solution through all its steps.
 Example: How many people live within 1 km from a railway in Bangladesh?”.
-In this group, 23 of the tasks were not solvable.
-Heatmaps, contour lines
-(54 tasks): required performing spatial operations (selection, overlaps, etc.), understanding the contents of rasters, and making a heatmap or generating contour lines as one of the steps.
+In this group, 22 of the tasks were not solvable.
+Heatmaps, contour lines (54 tasks): required performing spatial operations (selection, overlaps, etc.), understanding the contents of rasters, and making a heatmap or generating contour lines as one of the steps.
 Example: Create a heatmap of earthquake occurrences in regions with high population growth”.
 This was the most challenging set, as the agent had to figure out if the heatmaps and contour lines tools were actually able to solve the task and to make correct decisions on a complex set of arguments.
-In this group, 41 tasks ended up being unsolvable with available data/tools, which was a test for LLMs ability to recognize an unsolvable task and generate a rejection rather than a hallucinated answer.
-4.1 Tasks creation
-For each group, seed tasks were sourced from industry professionals and publications.
-Then more tasks were generated by using the Sonnet 3.5 model based on the seed tasks, the description of the tools, and the lists of available datasets.
-Then tasks were generated that should look similar to solvable tasks but could not be solved using the provided data and tools.
-All the generated tasks were then manually reviewed to keep only high-quality tasks in the set.
-4.2 Tasks annotation
+In this group, 27 tasks were not solvable.
+4.2. Tasks creation
+For each group, the seed tasks were sourced from industry professionals and publications.
+Then more tasks were generated by using the Claude 3.5 Sonnet model based on: the seed tasks, description of the tools, and the lists of available datasets.
+Unsolvable and solvable tasks were generated separately.
+The generated tasks were then manually reviewed to keep only high-quality tasks in the set.
+4.3. Tasks annotation
+To allow the evaluation of LLM-generated solutions, each task was annotated with the ground truth, reference solution, formatted as snippets of Python code (see more in Section 5).
+To create the annotation, we have started by generating draft reference solutions by the agent using GPT‑4o or Claude 3.5 Sonnet models.
+We then manually reviewed and rewrote every solution.
 Each of the 200+ tasks was annotated with one or multiple ground-truth reference solutions.
-To create the annotation, we have started by generating the draft reference solutions by the agent using GPT-4o or Sonnet 3.5 models.
-We then manually reviewed and rewrote every solution.
-One reference solution could be (1) Sequence of tool calls with arguments; similar to a Python code snippet (2) Empty sequence [] of tool calls meaning that the task can be solved from the LLM’s world knowledge (3) A single special tool call reject_task() meaning that the task should be rejected as it cannot be solved with given tools/data.
-A nonempty reference solution looked as follows for the evaluator:
+Reference solution could be
+  1. (1)
+    Sequence of tool-calls with arguments; similar to a python code snippet
+  2. (2)
+    Empty sequence [] of tool-calls meaning that the task can be solved from the LLM’s world knowledge
+  3. (3)
+    A single special tool call reject_task meaning that the task should be rejected as it cannot be solved with given tools/data
+A non-empty reference solution looked as follows for the evaluator:
 #overall reference solution comment
-#toolcall1 comment
-toolcall1(arg11=..., arg12=..., ...)
-#toolcall2 comment
-toolcall2(arg21=..., arg12=..., ...)
-Approximately 75% of the solvable tasks have reference solutions consisting of 4-6 steps (tool-calls) and  25% consist of more steps (up to 14 steps).
-We used comments in annotations as a way to record information important for the evaluator agent to make a judgment about whether the candidate solution and the reference solution are semantically equivalent.
-For example, telling which variations in the solution or step are irrelevant or important.
-One task sometimes allowed multiple solutions because tasks in practice are often ambiguous and multiple analytical approaches are possible for them.
-For example, task Compare the freshwater withdrawal between African countries with and without significant railway networks” can be solved: (1) by rejecting because it is not specific enough (user need to clarify what’s “signigicant” railway networks; (2) by making separate maps of railway networks and water withdrawals for African countries; (3) by combining the data and mapping bivariate map to show the relation; (4)by implying some threshold and computing statistical metrics for 2 groups of countries.
-5 Evaluation approach
-5.1 Matching-based evaluator
-The 2 widely used evaluation approaches are execution-based and matching-based.
-However, comparing directly the outputs of the agent on geospatial analytical tasks would be quite challenging because the outputs contain multiple images/maps, tables, numbers, a narrative and sometime all of it as in the section 4.2 example.
-Annotation of such ground truth is not practical in this case, so we opted for matching-based approach.
-We used LLM-as-Judge with multiple LLM judges to compare the solutions generated by the agent (“candidate solutions”) and reference solutions for the same task.
-To leverage the LLMs ability to analyze code, we represented both reference and candidate solutions as snippets of Python code as in 4.2.
-We asked LLM judges to compare semantic equivalence of the solution; this took care of situations where reference and candidate solutions had different order of steps, different arguments for tool-calls or even took a different path to the same analytical outcome.
-We also compared the lengths of the reference and candidate solutions, since the agent often takes some extra steps to figure out a path to answering the user question, and such extra steps are often costly.
-For evaluation of the solutions, we used a panel of judges: Sonnet 3.5, GPT-4o, and Gemini 2.0.
-The evaluator prompt included: persona setting, description of all tools, scoring taxonomy, instructions on how evaluate solutions, examples of evaluated solutions, corresponding scores, and reasoning behind each score.
-The evaluator was instructed to analyze the candidate solution for the task and produce (1) reasoning for matching any of the reference solutions for the task and the candidate solution and (2) the matching score (‘2’: match, ‘0’: no match, ‘1’: partial match)
-5.2 Evaluating the evaluator
-We evaluated and tuned the evaluator itself on a set of 44 tasks randomly sampled with stratification of all tasks, respecting task groupings.
-We manually scored candidate solutions for these tasks, providing a matching score and reasoning for it.
-We then ran the evaluator on the same tasks candidate solutions and compared human annotator match scores, and the evaluator produced match scores.
-Among the three models used as judges, the Sonnet 3.5 model matching scores were closest to human annotator scores.
-scoring (0.91 of the scores corresponded to manual scores), followed by GPT-4o (0.86) and Gemini 2.0 (0.82).
-In the below results section, we use all the 3 above LLMs as judges and average their match frequency estimates.
-6 LLMs evaluation results
-The following models were tested in the solver agent: gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18, o3-mini-2025-01-31, gemini-2.0-flas%h-001, claude-3-5-sonnet-20241022, claude-3-5-haiku-20241022, claude-3-7-sonne%t-20250219.
-Each LLM was tested on the same set of tasks, and all models were tested at temperature 0 (except o3-mini, which does not have a temperature parameter).
-The frequency of each of the match scores was calculated (1) separately for the solvable tasks (104 tasks) and the unsolvable tasks (98 tasks) (2) for each group of tasks according to 4.
-To mark unsolvable tasks, the agent was provided with a rejecttask tool that it could call either upfront or after some steps if it reached the conclusion that the task cannot be solved with available datasets and tools.
-To benchmark how efficient an LLM is at solving the task compared to the reference solution, we calculated the mean difference between the length (number of steps) of the reference solution (or average length of reference solutions if multiple) and the length of candidate solution.
-We express this difference relative to the length of the reference solution (so 0.5 means 50% more steps in the candidate solution than the reference).
-The first observation is that models that are more successful in identifying unsolvable tasks tend to be less successful in dealing with solvable tasks and vice versa (Table 1, Table 2).
-Anthropic models perform better on solvable tasks, but prioritize giving any answer at all over rejecting tasks, Sonnet 3.5 is the most balanced from the three in this aspect.
-Gemini 2.0 performs similarly on both kinds of tasks.
-Although the three OpenAI models are better at rejecting tasks than solving them, among them, the GPT-4o is well balanced.
-Overall, on solvable tasks, Sonnet 3.5 turned out to be the best model, followed by Sonnet 3.7, GPT-4o, Gemini 2.0 and Haiku 3.5 (Table 2).
-We can also see that for unsolvable tasks, the candidate solutions are much longer, indicating LLMs trying to find a solution and eventually giving up.
-For solvable tasks, candidate solutions are about the same length as reference ones, with notable exceptions of Sonnet 3.7 that tends to produce longer solutions than needed and o3-mini that tends to reject tasks more frequently than other models.
-The detailed statistics for each group of tasks are in tables 3, 4, 5, 6.
-Model No Partial Match $\Delta$ in
-match match solution
-length
-Sonnet 3.7 0.84 0.02 0.14 6.47
-Sonnet 3.5 0.61 0.02 0.37 2.51
-Haiku 3.5 0.79 0.01 0.20 4.39
-GPT-4o 0.38 0.03 0.59 2.45
-GPT-4o mini 0.37 0.02 0.62 2.71
-o3-mini 0.23 0.02 0.76 0.20
-Gemini 2.0 0.54 - 0.46 1.83
-Table 1: Success rate in identifying the unsolvable tasks (98 tasks from total of 202)
-Model No Partial Match $\Delta$ in
-match match solution
-length
-Sonnet 3.7 0.27 0.23 0.50 0.39
-Sonnet 3.5 0.28 0.19 0.53 -0.05
-Haiku 3.5 0.34 0.23 0.43 0.03
-GPT-4o 0.41 0.15 0.44 -0.01
-GPT-4o mini 0.70 0.13 0.16 -0.03
-o3-mini 0.69 0.10 0.21 -0.59
-Gemini 2.0 0.43 0.14 0.43 -0.03
-Table 2: Success rate in solvable tasks (104 tasks from total of 202)
-Model No Partial Match $\Delta$ in
-match match solutions
-length
-Sonnet 3.7 0.23 0.11 0.66 0.91
-Sonnet 3.5 0.17 0.15 0.72 0.29
-Haiku 3.5 0.25 0.21 0.54 0.51
-GPT-4o 0.24 0.13 0.63 0.31
-GPT-4o mini 0.49 0.19 0.32 0.31
-o3-mini 0.50 0.14 0.36 -0.39
-Gemini 2.0 0.25 0.06 0.68 0.16
-Table 3: Success rate in solving ‘Merge-visualize’ tasks (36 tasks from total of 202)
-Model No Partial Match $\Delta$ in
-match match solution
-length
-Sonnet 3.7 0.50 0.14 0.36 2.37
-Sonnet 3.5 0.30 0.10 0.60 0.40
-Haiku 3.5 0.59 0.09 0.33 2.11
-GPT-4o 0.26 0.09 0.65 0.29
-GPT-4o mini 0.54 0.07 0.39 0.60
-o3-mini 0.48 0.05 0.46 -0.38
-Gemini 2.0 0.35 0.12 0.53 0.17
-Table 4: Success rate in solving ‘Process-merge-visualize’ tasks (56 tasks from total of 202)
-Model No Partial Match $\Delta$ in
-match match solution
-length
-Sonnet 3.7 0.55 0.14 0.31 3.42
-Sonnet 3.5 0.48 0.12 0.40 1.73
-Haiku 3.5 0.50 0.17 0.33 2.49
-GPT-4o 0.50 0.08 0.42 1.95
-GPT-4o mini 0.69 0.05 0.26 2.44
-o3-mini 0.55 0.07 0.39 -0.3
-Gemini 2.0 0.63 0.07 0.30 2.07
-Table 5: Success rate in solving ‘Spatial operations’ tasks (53 tasks from total of 202)
-Model No Partial Match $\Delta$ in
-match match solution
-length
-Sonnet 3.7 0.81 0.13 0.06 6.05
-Sonnet 3.5 0.76 0.10 0.13 2.15
-Haiku 3.5 0.82 0.06 0.12 3.06
-GPT-4o 0.56 0.08 0.36 2.00
-GPT-4o mini 0.44 0.03 0.53 1.62
-o3-mini 0.39 - 0.61 0.17
-Gemini 2.0 0.64 0.04 0.32 0.91
-Table 6: Success rate in solving ‘Heatmaps, contour lines’ tasks (54 tasks from total of 202)
+toolcall1(arg11=..., arg12=..., ...) #call1 comment
+toolcall2(arg21=..., arg12=..., ...) #call2 comment
+One task sometimes allows multiple solutions because tasks in practice are often ambiguous and multiple analytical approaches are possible for them based on the provided data and tools.
+For example, task Compare the freshwater withdrawal between African countries with and without significant railway networks” can be solved:
+  • by rejecting because it’s not specific enough (user needs to clarify what’s “significant” railway networks);
+  • by making separate maps of railway networks and water withdrawals for African countries;
+  • by combining the data and mapping bivariate map to show the relation;
+  • by implying some threshold and computing statistical metrics for 2 groups of countries.
+We also used comments as a way to record some information important for the evaluator to make a judgment about whether the candidate solution and the reference solution are semantically equivalent.
+For example, providing alternative arguments to the tool call, specifying variations in arguments or marking optional steps.
+Approximately 74% of the solvable tasks have reference solutions consisting of 4-6 steps (tool-calls) and  28% consist of more steps (up to 14 steps).
+The numbers do not total to 100% as a task can have multiple reference solutions.
+5. Evaluation approach
+5.1. Matching-based evaluator
+To leverage the LLMs ability to comprehend code (Nikiema et al., 2025; Haroon et al., 2025; Ma et al., 2024), we represented both reference and candidate solutions as snippets of Python code as in 4.3.
+In code evaluation, the two widely used approaches are: execution-based and matching-based, both used for the evaluation of LLM-generated code (Wang and Chen, 2023; Anjum Haque et al., 2023).
+Since the outputs of agent on geospatial analytical tasks contain multiple images/maps, numbers, a narrative, and sometimes all of it as in the section 4.3 example, reliably comparing them via execution-based approach would be challenging and annotation of tasks with such ground truth is rarely practical.
+We opt for the matching-based approach for practicality and efficiency.
+We used reference-based LLM-as-Judge with multiple LLM judges to compare the solutions generated by the agent (“candidate solutions”) and reference solutions for the same task.
+Compared to human evaluation, the use of LLMs to compare the generated solution and reference is cost- and time-efficient (Chiang and Lee, 2023); if multiple rounds of evaluation are needed due to changes in tasks, their number, or evaluated LLMs, it offers reproducibility (Chiang and Lee, 2023) and scalability (Zheng et al., 2023).
+We asked LLM judge to compare semantic equivalence of the solution; this took care of situations where reference and candidate solutions had different order of steps, different arguments for tool-calls or even took a different path to the same analytical outcome.
+We also compared the lengths of the reference and candidate solutions, since the agent often takes some steps to figure out a path to answering the user question, and such extra steps might be costly (see 6.4).
+Looking into solution code also provides interesting insights into how LLM solves the tasks and allows agent tuning (which we did not do for this benchmarking).
+For evaluation of the solutions we used a panel of judges: Claude Sonnet 3.5, GPT-4.1 and Gemini 2.5 Pro Preview.
+The evaluator prompt included: persona setting, description of all tools, scoring taxonomy, instructions on how evaluate solutions, examples of evaluated solutions, corresponding scores and reasoning behind each score.
+The evaluator was instructed to analyze the candidate solution for the task and produce (1) reasoning for matching any of the reference solutions for the task and the candidate solution and (2) the matching score (’2’: match, ’0’: not a match, ’1’: partial match).
+5.2. Evaluating the evaluator
+We evaluated and tuned the evaluator itself on the set of 50 tasks (about 25% of the benchmark set) randomly sampled with stratification from all tasks, respecting the task groups.
+We manually scored candidate solutions for these tasks, providing a match score and reasoning for it.
+We then ran the evaluator on the same tasks candidate solutions and we compared human annotator match scores and evaluator produced match scores.
+Nine models were tested as judges (Gemini 2.0 Flash, Gemini 2.5 Pro Preview, Gemini 2.5 Flash Preview; Claude Sonnet 3.5, 3.7, 4; GPT-4o, GPT-4.1, GPT-4.1 mini) and from each family of models the one with the results closest to human annotator scores was selected for the panel: GPT-4.1 with 96% of scores matching the manual scores, Claude Sonnet 3.5 (94%) and Gemini 2.5 Pro Preview (88%) (see Figure 1).
+The top models of each family were selected rather than simply the top three models to reduce variance in the evaluation results due to differences in training and tuning of the models by different providers.
+Figure 1. Testing of LLMs as judges:
+In the evaluator, we use all the 3 above LLMs as judges and average their match frequency estimates.
+6. LLMs evaluation results
+6.1. Experiment procedure
+Nine models were tested in the solver agent: gemini-2.0-flash-001, gemini-2.5-pro-preview-05-06, claude-3-5-haiku-20241022, claude-3-5-sonnet-20241022, claude-sonnet-4-20250514, gpt-4o-2024-08-06, gpt-4.1-2025-04-14, o3-mini-2025-01-31, o4-mini-2025-04-16.
+Each LLM was tested on the same benchmark set of tasks, at temperature 0 (except o3-mini, o4-mini which do not support the temperature parameter).
+The frequency of each of the match scores was calculated:
+  1. (1)
+    separately by each group of tasks as per 4.1 and
+  2. (2)
+    for solvable tasks (123 tasks) and unsolvable tasks (79 tasks) across all groups.
+To benchmark how efficient an LLM is at solving a task compared to the reference solution, we also calculated the mean difference between the length (number of steps) of the reference solution for the task (or average length of reference solutions if multiple) and the candidate solution.
+We express this difference relative to the length of candidate solution (so 0.5 means 50% more steps in the candidate solution).
+The o3-mini model generated empty solutions (no tool calls, just thinking) for 103 tasks and rejected 75 tasks, and while the accuracy in identification of unsolvable tasks thus was 0.75, the model was excluded from further analysis due to obvious inefficiency in using tools to solve tasks.
+6.2. Performance on solvable and unsolvable tasks
+The first observation is that some models demonstrate a trade-off between the ability to identify and reject unsolvable tasks and the ability to solve tasks (Table 1, Table 2, Table 7, Figure 2).
+Claude Sonnet 4 and Haiku 3.5 perform better on solvable tasks but prioritize giving any answer over rejecting tasks, with Claude Sonnet 4 being a prominent case with a success rate on solvable tasks more than three times higher than on unsolvable ones.
+Gemini 2.5 Pro Preview is somewhat better at detecting unsolvable tasks than solving solvable ones.
+GPT-4o and o4-mini are substantially better in detecting unsolvable tasks than in solving feasible ones.
+In each family of models, there is one that demonstrated similar accuracy on both types of tasks in the experiment: Claude Sonnet 3.5, Gemini 2.0 Flash and GPT-4.1.
+We can also see, that for unsolvable tasks, the candidate solutions are much longer indicating LLMs trying to find a solution and eventually giving up.
+For solvable tasks, candidate solutions are about the same length as reference ones, with the exception of Claude Sonnet 4 which tends to produce longer solutions than needed.
+Overall, in solvable tasks, Claude Sonnet 4 turned out to be the best model, followed by o4-mini, Claude Sonnet 3.5, GPT-4.1, and the rest of the models were very close in their success rates (Table 2).
+However, the highest success rate is only slightly more than half (0.55 and 0.51), and most of the tested models are successful in less than half of the cases (note again that we specifically chose a trivial agent architecture to benchmark the ability of LLMs themselves rather than a particular agentic architecture).
+Some models have a relatively high proportion of solutions that partially match ground truth solutions (0.16-0.22 for Claude Haiku 3.5, GPT-4.1, Claude Sonnet 4, and Gemini 2.5 Pro Preview), meaning these solutions are very close to correct and models success rates might be increased with minor changes to the agent and tools.
+Table 1. Success rate in identifying the unsolvable tasks
+(79 tasks from total of 202) Model No Partial Match $\Delta$ in match match solution length Gemini 2.0 Flash 0.60 0.00 0.40 2.3 Gemini 2.5 Pro Preview 0.45 0.00 0.55 2.0 Claude Haiku 3.5 0.78 0.00 0.22 4.2 Claude Sonnet 3.5 0.53 0.00 0.47 2.5 Claude Sonnet 4 0.83 0.00 0.17 7.6 GPT-4o 0.38 0.00 0.62 2.5 GPT-4.1 0.57 0.00 0.43 4.5 o3-mini 0.25 0.00 0.75 0.0 o4-mini 0.10 0.00 0.90 0.5
+Table 2. Success rate in solvable tasks
+(123 tasks from total of 202) Model No Partial Match $\Delta$ in match match solution length Gemini 2.0 Flash 0.52 0.08 0.39 0.0 Gemini 2.5 Pro Preview 0.44 0.16 0.39 0.0 Claude Haiku 3.5 0.38 0.22 0.40 0.2 Claude Sonnet 3.5 0.41 0.12 0.47 -0.1 Claude Sonnet 4 0.27 0.18 0.55 0.7 GPT-4o 0.51 0.12 0.38 0.1 GPT-4.1 0.38 0.20 0.42 0.4 o4-mini 0.38 0.11 0.51 0.1
+Figure 2. Success rate (share of matches with reference solutions) across LLMs and task types:
+6.3. Performance by task groups
+We then examine LLMs performance by task groups without distinguishing between solvable and unsolvable tasks within each group.
+In the ’Merge - visualize’ task group (Table 3), the easiest and, mostly, solvable, almost all models achieved accuracy over 0.6.
+The o4-mini had the highest accuracy, followed by Claude Sonnet 3.5, 4, all with success rates over 0.7.
+Table 3. Success rate in solving ’Merge - visualize’ tasks
+(36 tasks from total of 202) Model No Partial Match $\Delta$ in match match solution length Gemini 2.0 Flash 0.32 0.00 0.68 0.2 Gemini 2.5 Pro Preview 0.20 0.13 0.67 0.7 Claude Haiku 3.5 0.25 0.15 0.60 0.6 Claude Sonnet 3.5 0.17 0.11 0.72 0.3 Claude Sonnet 4 0.19 0.09 0.71 1.3 GPT-4o 0.32 0.08 0.60 0.4 GPT-4.1 0.28 0.12 0.60 0.8 o4-mini 0.20 0.03 0.77 0.1
+In the ’Process - merge - visualize’ task group, accuracy starts to decrease while the solution lengths increase (Table 4).
+The top three models by accuracy achieve success rates over 0.6: GPT-4o, Claude Sonnet 3.5, and o4-mini, with o4-mini notably outperforming the other eight models at 0.75 accuracy.
+The rest have accuracy 0.5 or lower.
+In this task group, the tendency of Claude Sonnet 4 to produce lengthy tasks becomes apparent: its solutions on average are 330% longer than the reference solutions.
+Claude Haiku 3.5, the next model by mean difference in length with reference solutions, has the solutions 190% longer on average.
+The longer solutions usually mean repeated or failed steps, unnecessary or failed visualizations or unnecessary data loads early in the solution.
+This is associated with more errors, with the final response sometimes being unusable even if it contains the answer, or with inability to complete solution within the recursion limit.
+It also confuses evaluators that tend to penalize excessively long solutions despite the instructions not to.
+Most common problems for these two models were attempts to provide solutions for unsolvable tasks, errors in data filtering, selection of datasets and dates for mapping, lack of awareness of recent updates (like expansion of BRICS) even for models released after the events.
+Table 4. Success rate in solving ’Process - merge - visualize’ tasks
+(56 tasks from total of 202) Model No Partial Match $\Delta$ in match match solution length Gemini 2.0 Flash 0.46 0.06 0.48 0.04 Gemini 2.5 Pro Preview 0.45 0.05 0.50 1.0 Claude Haiku 3.5 0.54 0.09 0.36 1.9 Claude Sonnet 3.5 0.36 0.00 0.61 0.3 Claude Sonnet 4 0.50 0.08 0.41 3.3 GPT-4o 0.32 0.07 0.61 0.2 GPT-4.1 0.44 0.06 0.51 1.3 o4-mini 0.23 0.03 0.75 0.4
+In the third task group, ’Spatial operations’ only o4-mini achieves a success rate 0.56 (Table 5).
+However, in this task group o4-mini rejected 62% of tasks, which is higher than for the whole set (49%), hence the shorter mean difference between length of the ground truth and o4-mini solutions.
+Claude Sonnet 3.5, Gemini 2.5 Pro Preview, GPT-4o follow with the success rates 0.38-0.40.
+Six of the tested models have solutions on average more than twice as long as the reference solutions and the issues associated with this.
+Table 5. Success rate in solving ’Spatial operations’ tasks
+(53 tasks from total of 202) Model No Partial Match $\Delta$ in match match solution length Gemini 2.0 Flash 0.65 0.10 0.26 1.9 Gemini 2.5 Pro Preview 0.46 0.16 0.39 0.7 Claude Haiku 3.5 0.56 0.18 0.26 2.4 Claude Sonnet 3.5 0.50 0.11 0.40 1.6 Claude Sonnet 4 0.54 0.12 0.34 3.8 GPT-4o 0.55 0.06 0.38 2.3 GPT-4.1 0.46 0.19 0.35 2.9 o4-mini 0.33 0.10 0.56 -0.08
+In the fourth task group, ’Spatial Operations’, the accuracy (success rate) for most of models is the lowest across all task groups (Table 6).
+The top three models in this task group are: o4-mini with an accuracy rate of 0.60, due to the better ability to identify unsolvable tasks, Gemini 2.5 Pro Preview, and GPT-4o.
+Table 6. Success rate in solving ’Heatmaps, contour lines’ tasks
+(54 tasks from total of 202) Model No Partial Match $\Delta$ in match match solution length Gemini 2.0 Flash 0.74 0.05 0.21 1.3 Gemini 2.5 Pro Preview 0.60 0.08 0.31 0.8 Claude Haiku 3.5 0.72 0.13 0.15 2.0 Claude Sonnet 3.5 0.73 0.07 0.20 1.4 Claude Sonnet 4 0.65 0.15 0.20 4.6 GPT-4o 0.63 0.07 0.30 1.2 GPT-4.1 0.61 0.13 0.25 2.8 o4-mini 0.32 0.09 0.60 0.5
+Table 7. Share of solutions matching across models and tasks groups
+Model Unsolvable Solvable Merge - Process - Spatial Heatmaps,
+tasks tasks visualize merge - operations contour
+visualize lines
+Gemini 2.0 Flash 0.40 0.39 0.68 0.48 0.26 0.21
+Gemini 2.5 Pro Preview 0.55 0.39 0.67 0.50 0.39 0.31
+Claude Haiku 3.5 0.22 0.40 0.60 0.36 0.26 0.15
+Claude Sonnet 3.5 0.47 0.47 0.72 0.61 0.40 0.20
+Claude Sonnet 4 0.17 0.55 0.71 0.41 0.34 0.20
+GPT-4o 0.62 0.38 0.60 0.61 0.38 0.30
+GPT-4.1 0.43 0.42 0.60 0.51 0.35 0.25
+o4-mini 0.90 0.51 0.77 0.75 0.56 0.60
+6.4. Token usage
+We also compared the number of tokens used by the models to solve the benchmark set (Table 8) since it influences costs and costs are a critical parameter in practice.
+Note that because solving the tasks is an iterative process and at every iteration the previous model responses are fed back into model, the number of generated output tokens correlates with measured input tokens.
+What is clear is that an increase in token usage and generation does not necessarily translate into accuracy in solving tasks.
+Anthropic models were on the higher end of the input/output tokens used.
+Claude Sonnet 4, due to its reasoning ability, creates 2-3 times more input tokens and up to 5 times more output tokens than competitors, yet achieves similar or worse accuracy than models like Claude Sonnet 3.5, GPT-4o, GPT-4.1 and Gemini 2.5 Pro Preview (Table 7).
+Anthropic models are more expensive in use: the most affordable of the three, Claude 3.5 Haiku, due to the amount of tokens generated ended up costing more than any OpenAI or Google models tested in this experiment to solve the benchmark set.
+Notably, the best performed on this benchmark set OpenAI’s o4-mini model was among the three models with the lowest amount of input tokens used and the highest amount output token (3 times of the Claude Sonnet 4 output and 8 times GPT-41, next OpenAI’s model by output).
+Table 8. Tokens used for generation of solutions for the whole benchmark set (single run)
 Model Input Output
 tokens (mln) tokens (k)
-Sonnet 3.7 24 269
-Sonnet 3.5 12 157
-Haiku 3.5 15 203
-GPT-4o 7.5 57
-GPT-4o mini 8 55
-o3-mini 3.5 561
-Gemini 2.0 7 40
-Table 7: Tokens used for generation of solutions for the whole set
-6.1 Error Analysis
+Gemini 2.0 Flash 7.5 40.7
+Gemini 2.5 Pro Preview 10.4 85.4
+Claude Haiku 3.5 14.9 196.0
+Claude Sonnet 3.5 12.2 151.2
+Claude Sonnet 4 26.9 291.6
+GPT-4o 8.0 61.2
+GPT-4.1 12.7 121.3
+o4-mini 7.9 956.6
+6.5. Error analysis
 Looking at cases where candidate solutions did not match reference, the main causes/types of errors were:
-(1) Poor understanding of geometrical relations: mistaking centroids for full polygons, attempting to get centroids of lines, and not understanding that centroids can be outside object boundaries, etc, etc.
-(2) Outdated world knowledge – good for established international groupings but fails with recent changes (e.g., BRICS expansion in 2024) or ambiguous groupings (Upper/Lower Nile countries).
-(3) Over-reliance on world knowledge - using internally known spelling of names rather than dataset values for filtering, despite having tools to check column values.
-(4) Inefficient data manipulation – merging without pre-filtering (e.g., merging with all USA counties instead of first filtering to specific state).
-(5) Skipping geographic filtering steps – keeping continent-level data when only country-level was needed.
-(6) Attempting to solve with partial data, trying to be helpful by using whatever geography was available instead of rejecting tasks with insufficient data coverage.
-(7) Modifying tasks to make them solvable rather than solving the original problem.
-7 Conclusions
-We have created a benchmark dataset for multistep geo-spatial task completion, introduced an evaluation framework, and evaluated the leading mainstream commercial LLMs on these tasks.
-The task profile and the evaluation approach reflect the practical requirements of GIS professionals and the use of LLM in a GIS-assistant mode.
-Experiments indicate that commercial LLMs with function calling can successfully solve geospatial tasks even with casual, ambiguous instructions.
-Performance varies by task complexity, spatial thinking requirements, and model helpfulness bias.
-Sonnet 3.5 and GPT-4o showed the best overall results and a balance between identifying unsolvable tasks and solving solvable ones.
-Anthropic models performed well on easier tasks, but their preference for providing solutions rather than rejecting tasks, high token usage, and price may disadvantage them for geospatial tasks.
-o3-mini and GPT-4o-mini overall had lower success rates, especially in easy tasks, though they excel at identifying unsolvable tasks and are affordable.
-Limitations
-We did not tune prompts individually for each LLM; we evaluate only solution algorithm not the final response (for rejects and final numbers); the benchmark set size is moderate; we didn’t evaluate the variance of the responses (only the binomial confidence bounds).
-Appendix A Prompts for solver agent
-The system prompt is in Figure 1, the rules prompt is in Figure 2.
+  1. (1)
+    Poor understanding of geometrical relations: e.g., taking values from centroid point of a polygon instead of calculating average value over polygon, trying to use centroids of lines while looking for geometries overlaps, not understanding that object’s centroid can lie outside the boundary of the object, etc.
+  2. (2)
+    Lack of recent world knowledge: LLMs worked well with established international groupings, but struggled with ambiguous groupings (Upper/Lower Nile countries) or with more recent changes (e.g., BRICS expansion in 2024), even if the event happened before the model’s cut-off date.
+  3. (3)
+    Over-reliance on the world knowledge: using internally known spelling of names rather than dataset values for filtering, despite having tools to check column values.
+  4. (4)
+    Inefficient data manipulations: merging without prefiltering (e.g., merging with all USA counties instead of first filtering to specific state); filtering by countries when filtering by continent/region was possible, and/or repeating filtering steps sometimes confusing LLM and derailing solutions.
+  5. (5)
+    Skipping steps that would narrow down the geodatasets to the required geography. E.g keeping all the North American lakes when the question was about USA.
+  6. (6)
+    Attempting to solve with partial data, trying to be helpful by using whatever geography was available instead of rejecting tasks with insufficient data coverage.
+  7. (7)
+    Modifying tasks to make them solvable rather than solving the original problem.
+  8. (8)
+    Hallucinating unjustified data processing steps, rejecting solvable tasks in the middle of solution, using some numbers for thresholds or names of countries in filters without justification or connection with provided datasets.
+  9. (9)
+    Answering/rejecting the task without using the tools and dataset provided.
+Looking at models’ mistakes, it is likely that all models will perform much better if there is human-in-the-loop to disambiguate vague situations or provide step-by-step instructions on how to solve the task and more detailed explanations of the geographical scope of the tasks.
+It was also observed that making the tools more narrow specialized improved the accuracy of the agent.
+7. Conclusions
+We have created a benchmark dataset for multistep geospatial tasks, introduced an evaluation framework and evaluated the leading mainstream commercial LLMs on these tasks.
+The analysis indicates that the commercially available LLMs with function calling can be used successfully to solve geospatial tasks even when tasks are stated casually, with ambiguity, missing information, and minimal initial instructions.
+The performance of the models varies by the complexity of geospatial tasks, the degree of spatial thinking required to solve them, and how strongly the model is skewed to provide any answer at all rather than reject an unsolvable task.
+In terms of comparative performance (see Table 7 and Figure 2), the OpenAI’s o4-mini showed the best results overall.
+Anthropic’s Claude Sonnet 3.5 ranked second overall by consistently placing in the top three for accuracy across multiple categories.
+The three other models with top-3 accuracy in 2-3 categories showed clear trade-off between the ability to solve tasks and to identify unsolvable ones: Claude Sonnet 4 performs more than three times better on solvable tasks than on unsolvable ones thus losing in accuracy in separate task groups; GPT-4o and Gemini 2.5 Pro are better at correctly rejecting tasks than at solving them and thus have better accuracy on task groups.
+The remaining models, GPT-4.1 and Google’s Gemini 2.0 Flash, are more balanced in rejecting and solving tasks and have lower accuracy across task groups.
+Based on the cost of running the models, the best options for geospatial agents might be OpenAI’s o4-mini due to its combination of moderate costs and high accuracy, and Google’s Gemini 2.0 Flash as the most cost-efficient model while still providing decent results.
+Anthropic models performed well on easy tasks and solvable tasks in general; however, their preference for trying to provide any solution rather than rejecting, high token usage, and price might be disadvantages when using them for geospatial tasks.
+8. Limitations
+We did not tune prompts individually for each LLM; we did not provide human feedback to the model to guide solution; we evaluate only solution algorithm not the final response (for rejects and final numbers); the benchmark set size is moderate; we did not evaluate the variance of the responses (only the binomial confidence bounds).
+Appendix A: Prompts for solver agent
+System prompt is in Figure 3, rules prompt is in Figure 4.
 System Prompt You are a geographer who answers the questions with maps whenever possible.
@@
 Today is {current_date}.
-Figure 1: System prompt for solver agent.
+Figure 3. System prompt for solver agent.
 Rules Prompt While solving the tasks, please, follow the next rules: - If a task is not a geospatial task, if it does not require calling provided tools and use provided datasets to solve, proceed with reply, no explanations needed, no need to call reject_task’.
@@
 When the information is missing try to infer the reasonable values.
-Figure 2: Rules prompt for solver agent.
-Appendix B GeoSpatial Tools List
-loaddata
+Figure 4. Rules prompt for solver agent.
+Appendix B: GeoSpatial Tools List
+loaddata:
 Loads statistical data from the catalog into a Pandas DataFrame.
@@
     outputdataframename: Name for the new DataFrame
-loadgeodata
+    state: Current graph state containing (Geo)DataFrames and messages
+loadgeodata:
 Loads vector geospatial data from the catalog into a GeoPandas GeoDataFrame.
@@
     outputgeodataframename: Name for the new GeoDataFrame
-getrasterpath
+    state: Current graph state containing (Geo)DataFrames and messages
+getrasterpath:
 Constructs path to raster data from the catalog.
     rasterdataset: Name of the raster dataset to access
-getrasterdescription
+    state: Current graph state containing (Geo)DataFrames and messages
+getrasterdescription:
 Gets description of a raster dataset including metadata and statistics.
@@
     outputvariablename: Optional name for storing results
-analyzerasteroverlap
+    state: Current graph state containing (Geo)DataFrames and messages
+analyzerasteroverlap:
 Analyzes overlap between two rasters and calculates statistics.
@@
     plotresult: Whether to display visualization
-getvaluesfromrasterwithgeometries
+    state: Current graph state containing (Geo)DataFrames and messages
+getvaluesfromrasterwithgeometries:
 Masks a raster using vector geometries and calculates statistics.
@@
     plotresult: Whether to display visualization
-mergedataframes
+    state: Current graph state containing (Geo)DataFrames and messages
+mergedataframes:
 Merges statistical and geospatial DataFrames using key columns.
@@
     outputdataframename: Name for the merged DataFrame
-getuniquevalues
+    state: Current graph state containing (Geo)DataFrames and messages
+getuniquevalues:
 Gets unique values from a column in a DataFrame/GeoDataFrame.
@@
     column: Column name to get unique values from
-filtercategorical
+    state: Current graph state containing (Geo)DataFrames and messages
+filtercategorical:
 Filters DataFrame/GeoDataFrame by categorical values.
@@
     outputdataframename: Name for filtered DataFrame
-filternumerical
+    state: Current graph state containing (Geo)DataFrames and messages
+filternumerical:
 Filters DataFrame/GeoDataFrame using numerical conditions.
@@
     outputdataframename: Name for filtered DataFrame
-calculatecolumnstatistics
+calculatecolumnstatistics:
 Calculates summary statistics for a numerical column.
@@
     additionalquantiles: Additional quantiles to calculate
-createbuffer
+    state: Current graph state containing (Geo)DataFrames and messages
+createbuffer:
 Creates buffer zones around geometries in a GeoDataFrame.
@@
     outputgeodataframename: Name for result GeoDataFrame
-makechoroplethmap
+    state: Current graph state containing (Geo)DataFrames and messages
+makechoroplethmap:
 Creates a choropleth map visualization from a GeoDataFrame.
@@
     colormap: Color scheme for the map
-filterpointsbyrastervalues
+    state: Current graph state containing (Geo)DataFrames and messages
+filterpointsbyrastervalues:
 Samples raster values at point locations and filters by threshold.
@@
     plotresult: Whether to display visualization
-selectfeaturesbyspatialrelationship
+    state: Current graph state containing (Geo)DataFrames and messages
+selectfeaturesbyspatialrelationship:
 Selects features based on spatial relationships.
@@
     plotresult: Whether to display visualization
-calculatelinelengths
+    state: Current graph state containing (Geo)DataFrames and messages
+calculatelinelengths:
 Calculates lengths of line features in kilometers.
@@
     outputvariablename: Name for storing results
-calculatecolumns
+    state: Current graph state containing (Geo)DataFrames and messages
+calculatecolumns:
 Performs operations between columns of DataFrames.
@@
     outputdataframename: Name for result DataFrame
-scalecolumnbyvalue
+    state: Current graph state containing (Geo)DataFrames and messages
+scalecolumnbyvalue:
 Performs operations between a column and a value.
@@
     outputdataframename: Name for result DataFrame
-makeheatmap
+    state: Current graph state containing (Geo)DataFrames and messages
+makeheatmap:
 Creates an interactive heatmap from point data.
@@
     valuecolumn: Column for intensity values
-  • outputhtmlname: Optional name for HTML output
     centerlat/lon: Map center coordinates
@@
     width/height: Map dimensions
-visualizegeographies
+    state: Current graph state containing (Geo)DataFrames and messages
+visualizegeographies:
 Displays multiple GeoDataFrames as map layers.
@@
     addlegend: Whether to add a legend
-getcentroids
+    state: Current graph state containing (Geo)DataFrames and messages
+getcentroids:
 Calculates centroids of polygon features.
@@
     outputgeodataframename: Name for centroid GeoDataFrame
-plotcontourlines
-Creates contour lines from raster data.
-  • rasterpath: Path to the raster file
-  • outputgeodataframename: Name for contour GeoDataFrame
-  • interval: Interval between contour lines
-  • min/maxvalue: Range for contours
-  • plotresult: Whether to display plot
-  • title: Plot title
-generatecontoursdisplay
+    state: Current graph state containing (Geo)DataFrames and messages
+generatecontoursdisplay:
 Creates contour lines using GDAL algorithm.
@@
     plotresult: Whether to display visualization
-makebivariatemap
+    plottitle: Title for the plot
+    cleanupfiles: Whether to remove temporary files after processing
+    colormap: Matplotlib colormap name
+    figsize: Figure size as JSON array [width, height]
+    addcolorbar: Whether to add a colorbar to the plot
+    plotbackground: Whether to display raster as background
+    state: Current graph state containing (Geo)DataFrames and messages
+makebivariatemap:
 Creates a bivariate choropleth map showing relationship between two variables.
@@
     outputvariablename: Name for storing results
-rejecttask
+    state: Current graph state containing (Geo)DataFrames and messages
+rejecttask:
 Returns a standardized message when task cannot be solved.
+    state: Current graph state containing (Geo)DataFrames and messages